# Low-Rank Adaptation for Foundation Models: A Comprehensive Review

### v1

@@ Abstract
The rapid advancement of foundation models—large-scale neural networks trained on diverse, extensive datasets—has revolutionized artificial intelligence, enabling unprecedented advancements across domains such as natural language processing, computer vision, and scientific discovery. However, the substantial parameter count of these models, often reaching billions or trillions, poses significant challenges in adapting them to specific downstream tasks. Low-Rank Adaptation (LoRA) has emerged as a highly promising approach for mitigating these challenges, offering a parameter-efficient mechanism to fine-tune foundation models with minimal computational overhead. This survey provides the first comprehensive review of LoRA techniques beyond large Language Models to general foundation models, including recent techniques foundations, emerging frontiers and applications of low-rank adaptation across multiple domains. Finally, this survey discusses key challenges and future research directions in theoretical understanding, scalability, and robustness. This survey serves as a valuable resource for researchers and practitioners working with efficient foundation model adaptation.
@@
Foundation models are characterized by their unprecedented scale, with parameter counts reaching billions or even trillions, and exhibit emergent properties - capabilities that arise spontaneously without explicit training [1]. These architectures have become fundamental building blocks of modern AI systems, enabling breakthrough performance across diverse domains [1, 2] While these models exhibit broad capabilities, task-specific optimization through fine-tuning remains essential for enhancing generalization [15], promoting algorithmic fairness [16], enabling customization [17], and aligning with ethical and societal standards [18, 19]. However, their scale introduces significant computational challenges, particularly in the computational resources required for both training and fine-tuning [20]. Although traditional fine-tuning methods involving full parameters updates have demonstrated effectiveness across various tasks [21, 22], their computational demands often render them impractical for foundation models [23, 24].
@@
LoRA is grounded on two key insights that weight updates during fine-tuning often reside in a low-dimensional subspace [29, 30], and task-specific adaptations can be effectively captured using low-rank matrices [17]. By optimizing these low-rank matrices per task while freezing the original model parameters, LoRA achieves efficient adaptation and enables the composition of multiple task-specific adaptations without increasing inference latency [31, 17].
@@ 2 Basics
LoRA [17] constitutes a substantial advancement in parameter-efficient fine-tuning (PEFT). Although originally developed for LLMs, subsequent research has demonstrated its effectiveness across a diverse of foundation models.
@@
(1) Parameter Efficiency. LoRA introduces a minimal set of trainable parameters through low-rank decomposition, typically reducing the number of task-specific parameters by several orders of magnitude. This approach is particularly advantageous in resource-constrained environments and multi-task scenarios where multiple adaptations of a base model are required.
(2) Enhanced Training Efficiency. Unlike conventional full fine-tuning, which updates all model parameters, LoRA optimizes only the low-rank adaptation matrices. This approach substantially reduces computational costs and memory requirements, especially for models with billions of parameters. The reduced parameter space typically leads to faster convergence during training.
(3) None-latency Inference. LoRA does not introduce additional inference latency since the update matrix can be explicitly incorporated into the original frozen weights . This integration ensures that the adapted model maintains efficiency during deployment and inference.
@@
(5) Robust Knowledge Retention. By preserving the pre-trained weights, LoRA effectively mitigates catastrophic forgetting, a common challenge in conventional fine-tuning. This approach maintains the model’s foundational knowledge while acquiring task-specific capabilities.
@@ 3.1.1 Parameter Decomposition
(1) Update Matrix Decomposition. In update matrix decomposition approaches, two primary strategies have emerged: singular value decomposition (SVD) based method and tensor train (TT)-based decomposition.
@@
LoRETTA introduced two variants: LoRETTA and LoRETTA. LoRETTA employs tensorized adapters, inserting these lightweight modules after each attention and feed-forward sub-layer in the transformer blocks. LoRETTArep, on the other hand, reparameterizes the weight matrix with tensor factors, offering an even more compact PEFT approach. It updates the weights using two unbiased tensorized layers, further reducing the number of trainable parameters while maintaining comparable performance.
@@
| Asymmetric LoRA [46] | Theoretical design | Orthogonal projection | Random orthogonal with theoretical guarantees |
@@ 3.1.4 Parameter Quantization
During-finetuning quantization. During-finetuning quantization applies quantization both before and throughout the fine-tuning process. Methods like QA-LoRA [57] leverage group-wise quantization to adjust the precision dynamically during training, ensuring a more balanced interaction between low-rank updates and quantized weights.
Post-finetuning quantization. Post-finetuning quantization, such as in LQER [58], occurs after the fine-tuning is completed, focusing on quantization primarily for inference. LQER utilizes a low-rank SVD-based decomposition to minimize quantization errors, ensuring that the quantized weights closely match the original high-precision weights.
@@
| QDyLoRA | During FT | Pretrained | mixed | Rank sampling | Dynamic LoRA | Rank selection | FT | None |
@@
(2) Quantization Technique. Different approaches include uniform quantization, non-uniform quantization, and mixed-precision quantization have been proposed for LoRA.
Uniform Quantization. Uniform quantization assigns the same bit-width across all weights, regardless of their distribution. QA-LoRA [57] applies uniform quantization with group-wise refinement, which balances the precision trade-offs to optimize memory efficiency and adaptation. However, uniform quantization may struggle with non-uniformly distributed weights, which is where non-uniform quantization becomes more effective.
Mixed-precision Quantization. Mixed-precision quantization offers even more flexibility by dynamically adjusting the bit-width based on the weight matrix or layer. Methods such as LoftQ [56] and LQ-LoRA [59] leverage mixed precision to optimize quantization across different components of the model. LoftQ, for example, alternates between quantizing the residuals of the weight matrix and refining the low-rank components using SVD. By iteratively optimizing the low-rank parameters and adjusting the quantization levels, LoftQ is able to minimize quantization errors. LQ-LoRA expands upon this approach by employing integer linear programming to dynamically configure the bit-width for each weight matrix, and also introduces a data-aware mechanism that uses an approximation of the Fisher information matrix to inform the quantization process. This allows LQ-LoRA to achieve a more accurate decomposition of the weight matrices with minimal quantization-induced loss.
To summarize, pre-finetuning quantization approaches, such as QLoRA and LoftQ, generally provide greater memory savings by freezing the pretrained weights, while post-finetuning methods like LQER focus more on refining precision for inference. In terms of quantization techniques, non-uniform and mixed precision methods, as seen in QLoRA, LoftQ, and LQ-LoRA, demonstrate superior performance in low-bit scenarios by offering more flexible precision allocation based on weight distribution. The timing of quantization and the specific quantization technique both play critical roles in determining the balance between memory efficiency and model performance. Table III provides a comprehensive summary of the discussed quantization methods.
Overall, parameter efficiency enhancement techniques in LoRA has evolved through four complementary approaches: decomposition, pruning, freezing and sharing, and quantization, each offering unique trade-offs between model performance and resource utilization. Building upon these efficiency gains, we next explore how rank adaptation strategies further advance LoRA’s capabilities.
3.2 Ranking Adaptation
Rank is a crucial parameter in LoRA, directly impacting the model adaptability and number of trainable parameters. The original LoRA method employs a fixed low rank across all layers, which may not be optimal for different downstream tasks and model architectures. To address limitations, recent works have proposed various approaches to optimize rank allocation in LoRA, which can be broadly categorized into two main aspects: rank refinement and rank augmentation. Figure 5 presents an illustration of these two methods.
@@ 3.2.1 Rank Refinement
Rank refinement methods aim to adaptively select the rank of LoRA modules during the fine-tuning. The key insight is that different layers may require varying degrees of adaptation and thus benefit from different ranks. Rank refinement approaches can be grouped into three main types: adaptive allocation, heuristic strategies, and multi-rank training.
(1) Adaptive Allocation. Adaptive allocation methods dynamically adjust the ranks of LoRA modules during training based on importance metrics derived from the data or model parameters. AdaLoRA [36] introduces an adaptive mechanism for rank allocation by parameterizing LoRA updates using SVD. It dynamically prunes singular values based on their magnitudes, allowing each layer to have a tailored rank while maintaining a global parameter budget. Similarly, SoRA [43] employs a learnable gating mechanism to control the effective rank of each LoRA module. The gates are optimized using proximal gradient descent with regularization to promote sparsity. This approach enables the automatic discovery of suitable ranks for different layers, improving parameter efficiency without manual tuning.
(2) Heuristic Strategies. Heuristic strategies allocate ranks based on predefined rules, which can be from prior knowledge or empirical observations. PRILoRA [60] proposes a deterministic strategy where the rank of LoRA modules increases linearly from lower to higher layers. Motivated by the observation that higher layers often require more adaptation in transfer learning, this heuristic allocates higher ranks to upper layers.
@@ 3.2.2 Rank Augmentation
Rank augmentation methods aim to achieve high-rank model updates through sequences of low-rank modifications, bridging the performance gap between LoRA and full-parameter fine-tuning. These methods can be categorized into two types: matrix merging-based methods and matrix resampling-based methods.
@@
(1) Matrix merging-based methods. Matrix merging-based methods increase the rank by merging low-rank update matrices. The key idea is that the sum of multiple low-rank matrices can approximate a higher-rank matrix, thereby enhancing the ability to capture complex patterns without incurring substantial computational overhead.
ReLoRA [62] introduces an iterative training framework where low-rank LoRA modules are trained and periodically merged into the pre-trained model weights. By resetting the optimizer and initializing new LoRA modules after each merge, ReLoRA effectively increases the overall rank while maintaining memory efficiency.
COLA [63] proposes a similar iterative optimization strategy inspired by the Frank-Wolfe algorithm [64]. It iteratively trains LoRA modules and merges them into the model, incrementally building a higher-rank adaptation. Each new LoRA module minimizes the residual error from previous adaptations, enabling COLA to achieve high-rank expressiveness without increasing the computational cost per iteration.
@@
MELoRA [65] introduces a parallelization approach to rank augmentation. The core idea is to train multiple small LoRA modules concurrently and concatenate their outputs to form a higher-rank adaptation. By assembling a mini-ensemble of low-rank adapters, MELoRA constructs an equivalent block-diagonal matrix that collectively has a higher rank.
XGBLoRA [66] introduces the Gradient Boosting (GB) framework for LoRA. It assembles the final model by combining a sequence of Rank-1 boosters (LoRA adaptations), progressively refining the model’s predictions. With the GB principle of the weak learner (i.e., strong ensemble model from a set of weak predictors), XGBLoRA overcomes the dilemma between extreme low-rank adaption and effectiveness.
@@
(2) Resampling-based method. Matrix resampling-based methods achieve high-rank adaptations by dynamically resampling the projection matrices during training. The fundamental idea is to utilize time to accumulate the effect of high-rank updates while operating with low-rank matrices at each training step.
@@
FLoRA [67] reinterprets LoRA as a gradient compression and decompression mechanism. The motivate that FLoRA periodically resamples the projection matrices used in LoRA modules during training. By changing these matrices at scheduled intervals, the method ensures that different subspaces are explored over time, effectively accumulating a higher-rank adaptation.
@@
In summary, rank adaptation strategies enhance LoRA adaptability by tailoring the rank of adaptation matrices to better suit the requirements of different layers and tasks. Table IV presents a detailed summary about rank refinement and augmentation.
@@ 3.3 Training Process Improvements
Learning Rate. In standard LoRA fine-tuning, a uniform learning rate is typically applied to both low-rank matrices and . However, Hayou et al. [68] observe that this practice leads to suboptimal performance, especially as the model width increases. The issue lies in the updates to and contribute differently to the learning dynamics. To address this limitation, Hayou et al. [68] propose LoRA+, a method that assigns different learning rates to matrices and . Their theoretical analysis in the infinite-width limit reveals that for efficient learning, the magnitudes of feature updates from both and should be . This necessitates scaling the learning rates such that and , where denotes the model width. In practice, LoRA+ introduces a fixed ratio , allowing practitioners to tune a single learning rate while automatically adjusting the other.
@@
By adaptive learning rates to the distinct roles of LoRA matrices, mitigating overfitting through structured dropout, and preventing gradient collapse with rank-stabilized scaling, these methods enhance both the efficiency and effectiveness of LoRA fine-tuning. We next examine the theoretical foundations underlying LoRA’s performance.

### v2

@@ Abstract
The rapid advancement of foundation models—large-scale neural networks trained on diverse, extensive datasets—has revolutionized artificial intelligence, enabling unprecedented advancements across domains such as natural language processing, computer vision, and scientific discovery. However, the substantial parameter count of these models, often reaching billions or trillions, poses significant challenges in adapting them to specific downstream tasks. Low-Rank Adaptation (LoRA) has emerged as a highly promising approach for mitigating these challenges, offering a parameter-efficient mechanism to fine-tune foundation models with minimal computational overhead. This survey provides the first comprehensive review of LoRA techniques beyond large language models to general foundation models, including recent technical foundations, emerging frontiers, and applications of low-rank adaptation across multiple domains. Finally, this survey discusses key challenges and future research directions in theoretical understanding, scalability, and robustness. This survey serves as a valuable resource for researchers and practitioners working with efficient foundation model adaptation. Code is available at https://github.com/marlin-codes/awesome-lora-adapter
@@
Foundation models are characterized by their unprecedented scale, with parameter counts reaching billions or even trillions, and exhibit emergent properties, capabilities that arise spontaneously without explicit training [1]. These architectures have become fundamental building blocks of modern AI systems, enabling breakthrough performance across diverse domains [1, 2]. While these models exhibit broad capabilities, task-specific optimization through fine-tuning remains essential for enhancing generalization [15], promoting algorithmic fairness [16], enabling customization [17], and aligning with ethical and societal standards [18, 19]. However, their scale introduces significant computational challenges, particularly in the computational resources required for both training and fine-tuning [20]. Although traditional fine-tuning methods involving full parameter updates have demonstrated effectiveness across various tasks [21, 22], their computational demands often render them impractical for foundation models [23, 24].
@@
LoRA is grounded on two key insights: weight updates during fine-tuning often reside in a low-dimensional subspace [29, 30], and task-specific adaptations can be effectively captured using low-rank matrices [17]. By optimizing these low-rank matrices per task while freezing the original model parameters, LoRA achieves efficient adaptation and enables the composition of multiple task-specific adaptations without increasing inference latency [31, 17].
@@
This survey organizes existing LoRA research as illustrated in Figure 2. It also highlights the critical challenges and future research directions in Section 6, with Figure 3 presenting a detailed structure diagram of the survey. This survey offers a valuable resource for both researchers and practitioners in the field.
@@ 2 Basics
LoRA [17] constitutes a substantial advancement in parameter-efficient fine-tuning (PEFT). Although originally developed for LLMs, subsequent research has demonstrated its effectiveness across a diverse set of foundation models.
@@
(1) Parameter Efficiency. LoRA introduces a minimal set of trainable parameters through low-rank decomposition, typically reducing the number of task-specific parameters by several orders of magnitude. This is especially beneficial in resource-constrained and multi-task scenarios requiring multiple model adaptations.
(2) Enhanced Training Efficiency. Unlike conventional full fine-tuning, which updates all model parameters, LoRA optimizes only the low-rank matrices, substantially reducing computational costs and memory requirements, especially for models with billions of parameters. The reduced parameter space typically leads to faster convergence during training.
(3) No Additional Inference Latency. LoRA does not introduce extra inference latency because the update matrix can be explicitly incorporated into the original frozen weights . This integration ensures that the adapted model maintains efficiency during deployment and inference.
@@
(5) Robust Knowledge Retention. By preserving the pre-trained weights, LoRA effectively mitigates catastrophic forgetting, a common challenge in conventional fine-tuning. LoRA maintains the model’s foundational knowledge while acquiring task-specific capabilities.
@@ 3.1.1 Parameter Decomposition
(1) Update Matrix Decomposition. In update matrix decomposition approaches, two primary strategies have emerged: singular value decomposition (SVD)-based methods and tensor train (TT)-based decomposition.
@@
LoRETTA introduces two variants: LoRETTA and LoRETTA. LoRETTA employs tensorized adapters, inserting these lightweight modules after each attention and feed-forward sub-layer in the transformer blocks. LoRETTArep, on the other hand, reparameterizes the weight matrix with tensor factors, offering an even more compact PEFT approach. It updates the weights using two unbiased tensorized layers, further reducing the number of trainable parameters while maintaining comparable performance.
@@ 3.1.2 Parameter Pruning
| AsyLoRA [47] | Theoretical design | Orthogonal projection | Random orthogonal with theoretical guarantees |
@@ 3.1.4 Parameter Quantization
During-fine-tuning quantization. During-fine-tuning quantization applies quantization both before and throughout the fine-tuning process. Methods like QA-LoRA [57] leverage group-wise quantization to adjust the precision dynamically during training, ensuring a more balanced interaction between low-rank updates and quantized weights. L4Q [58] merges the LoRA parameters and with the original weight matrix to form a new matrix, which is then quantized, enabling simultaneous optimization of both the quantization and LoRA parameters.
Post-fine-tuning quantization. Post-fine-tuning quantization, such as in LQER [59], occurs after fine-tuning is completed, focusing on quantization primarily for inference. LQER utilizes a low-rank SVD-based decomposition to minimize quantization errors, ensuring that the quantized weights closely match the original high-precision weights.
(2) Quantization Techniques. Approaches for LoRA include uniform quantization, non-uniform quantization, and mixed-precision quantization.
Uniform Quantization. Uniform quantization assigns the same bit-width across all weights, regardless of their distribution. QA-LoRA [57] applies uniform quantization with group-wise refinement, balancing the precision trade-offs to optimize memory efficiency and adaptation. Remarkably, LowRA [60] utilizes a data-free post-training quantization approach, reducing the quantization precision to below 2 bits through low-rank tensor optimization. However, uniform quantization may struggle with non-uniformly distributed weights, which is where non-uniform quantization becomes more effective.
@@
| L4Q [58] | Pre & During FT | Pretrained | 2, 3, 4 bit | Uniform & NormalFloat | Q-aware LoRA | Joint | FT&Inference | None |
| LowRA [60] | Pre-FT | Pretrained | 2 bit | Uniform & Group-wise | Standard LoRA | Separate | FT | None |
@@
| QDyLoRA [61] | During FT | Pretrained | mixed | Rank sampling | Dynamic LoRA | Rank selection | FT | None |
@@
Mixed-precision Quantization. Mixed-precision quantization provides more flexibility by dynamically adjusting the bit-width based on the weight matrix or layer. Methods like LoftQ [56] and LQ-LoRA [62] optimize quantization across model components. LoftQ quantizes the residuals of the weight matrix and refines low-rank components using SVD, iteratively minimizing quantization errors. LQ-LoRA extends this approach by using integer linear programming to dynamically configure the bit-width for each weight matrix and incorporates a data-aware mechanism, utilizing an approximation of the Fisher information matrix to improve accuracy while minimizing quantization loss.
To summarize, pre-fine-tuning quantization approaches like QLoRA and LoftQ generally provide greater memory savings by freezing the pretrained weights, while post-fine-tuning methods like LQER focus more on refining precision for inference. In terms of quantization techniques, non-uniform and mixed-precision methods, as seen in QLoRA, LoftQ, and LQ-LoRA, demonstrate superior performance in low-bit scenarios by offering more flexible precision allocation based on weight distribution. The timing of quantization and the specific quantization technique both play critical roles in determining the balance between memory efficiency and model performance. Table III provides a comprehensive summary of the discussed quantization methods.
Overall, parameter efficiency enhancement techniques in LoRA have evolved through four complementary approaches: decomposition, pruning, freezing and sharing, and quantization, each offering unique trade-offs between model performance and resource utilization. Building upon these efficiency gains, we next explore how rank adaptation strategies further advance LoRA’s capabilities.
3.2 Rank Adaptation
Rank is a crucial parameter in LoRA, directly impacting the model adaptability and number of trainable parameters. The original LoRA method employs a fixed low rank across all layers, which may not be optimal for different downstream tasks and model architectures. To address this limitation, recent works have proposed various approaches to optimize rank allocation in LoRA, which can be broadly categorized into two main aspects: rank refinement and rank augmentation. Figure 5 presents an illustration of these two methods.
@@ 3.2.1 Rank Refinement
Rank refinement methods aim to adaptively select the rank of LoRA modules during fine-tuning. The key insight is that different layers may require varying degrees of adaptation and thus benefit from different ranks. Rank refinement approaches can be grouped into three main types: adaptive allocation, heuristic strategies, and multi-rank training.
(1) Adaptive Allocation. Adaptive allocation methods dynamically adjust the ranks of LoRA modules during training based on importance metrics derived from the data or model parameters. AdaLoRA [36] introduces an adaptive mechanism for rank allocation by parameterizing LoRA updates using SVD. It dynamically prunes singular values based on their magnitudes, allowing each layer to have a tailored rank while maintaining a global parameter budget.
Similarly, SoRA [44] employs a learnable gating mechanism to control the effective rank of each LoRA module. The gates are optimized using proximal gradient descent with regularization to promote sparsity. This approach enables the automatic discovery of suitable ranks for different layers, improving parameter efficiency without manual tuning.
(2) Heuristic Strategies. Heuristic strategies allocate ranks based on predefined rules derived from prior knowledge or empirical observations. PRILoRA [63] proposes a deterministic strategy where the rank of LoRA modules increases linearly from lower to higher layers. Motivated by the observation that higher layers often require more adaptation in transfer learning, this heuristic allocates higher ranks to upper layers.
@@ 3.2.2 Rank Augmentation
Rank augmentation methods aim to achieve high-rank model updates through sequences of low-rank modifications, bridging the performance gap between LoRA and full-parameter fine-tuning. These methods can be categorized into three types: matrix merging-based, high-rank updating-based, and matrix resampling-based.
@@
(1) Matrix merging-based methods increase the rank by merging low-rank update matrices. The key idea is that the sum of multiple low-rank matrices can approximate a higher-rank matrix, thereby enhancing the ability to capture complex patterns without incurring substantial computational overhead.
ReLoRA [65] and COLA [66] introduce iterative optimization strategies to increase model rank. Both methods focus on periodically merging low-rank modules into pre-trained weights. ReLoRA employs an iterative framework where LoRA modules are periodically merged, resetting the optimizer and training new LoRA modules after each merge. Similarly, COLA, inspired by the Frank-Wolfe algorithm [67], merges LoRA modules in a sequence, optimizing them to reduce residual errors and achieve high-rank expressiveness.
@@
MELoRA [68] introduces a parallelization approach to rank augmentation, which trains multiple small LoRA modules concurrently and concatenates their outputs to form a higher-rank adaptation. By assembling a mini-ensemble of low-rank adapters, MELoRA constructs an equivalent block-diagonal matrix that collectively has a higher rank.
XGBLoRA [69] integrates the Gradient Boosting (GB) framework into LoRA by combining Rank-1 boosters in sequence to refine the model’s predictions, overcoming the dilemma between extremely low-rank adaptation and effectiveness.
LoRA-LEGO [70] performs rank-wise parameter clustering, grouping Minimal Semantic Units (MSUs) from different LoRAs into clusters and optimizes their scale using a dual reweighting strategy, achieving rank augmentation by merging low-rank matrices while maintaining computational efficiency.
(2) High-Rank Updating-Based methods enhance model capacity by increasing the rank of weight updates, typically using low-rank or non-trainable components, while maintaining computational and parameter efficiency.
MoRA [71] introduces a method for high-rank updating by using a square matrix, achieving a higher rank in the weight updates while maintaining the same number of trainable parameters. In addition, the method utilizes non-parameterized operators to reduce the input dimension and increase the output dimension for the square matrix, ensuring that the weight can still be merged back into the model.
HiRA [72] utilizes the Hadamard product to construct a high-rank update matrix. Additionally, the authors propose HiLoRA, a combination of HiRA and LoRA, demonstrating that high-rank updates can provide greater capacity without additional inference overhead.
RandLoRA [73] constructs a high-rank update matrix by learning linear combinations of low-rank, non-trainable random matrices. During training, only the scaling coefficients of these matrices are optimized, enabling full-rank updates while preserving parameter efficiency. Additionally, it leverages sparse random matrices to further improve memory and computational efficiency, inspired by random projection techniques.
@@
(3) Resampling-based methods achieve high-rank adaptations by dynamically resampling projection matrices during training, accumulating the effect of high-rank updates while using low-rank matrices at each step.
@@
FLoRA [74] reinterprets LoRA as a gradient compression and decompression mechanism. It periodically resamples the projection matrices used in LoRA modules during training. By changing these matrices at scheduled intervals, the method ensures that different subspaces are explored, effectively accumulating a higher-rank adaptation.
@@
| LoRA-LEGO[70] | Augmentation | Matrix Merging-Based |
| MoRA[71] | Augmentation | High-Rank Updating |
| HiRA[72] | Augmentation | High-Rank Updating |
| RandLoRA[73] | Augmentation | High-Rank Updating |
@@
In summary, rank adaptation strategies enhance LoRA adaptability by tailoring the rank of adaptation matrices to better suit the requirements of different layers and tasks. Table IV presents a detailed summary of rank refinement and augmentation.
@@ 3.3 Training Process Improvements
Learning Rate. In standard LoRA fine-tuning, a uniform learning rate is typically applied to both low-rank matrices and . However, Hayou et al. [75] observe that this practice leads to suboptimal performance, especially as the model width increases. The issue lies in the fact that updates to and contribute differently to the learning dynamics. To address this limitation, Hayou et al. [75] propose LoRA+, a method that assigns different learning rates to matrices and . Their theoretical analysis in the infinite-width limit reveals that for efficient learning, the magnitudes of feature updates from both and should be . This necessitates scaling the learning rates such that and , where denotes the model width. In practice, LoRA+ introduces a fixed ratio , allowing practitioners to tune a single learning rate while automatically adjusting the other.
@@
By adapting learning rates to the distinct roles of LoRA matrices, mitigating overfitting through structured dropout, and preventing gradient collapse with rank-stabilized scaling, these methods enhance both the efficiency and effectiveness of LoRA fine-tuning. We next examine the theoretical foundations underlying LoRA’s performance.